# MOSS: Motion-based 3D Clothed Human Synthesis from Monocular Video

### v1

@@ 2.4 Surface Deformation Detector
We densify the Gaussians based on gradient localization to increase the rendering granularity for better reproduction of clothing folds and joint details. Remember that the set of Gaussians selected by the gradient is $M=\{m1,m2,...\}$. For each point $m_{i}$ in the point set $M$, select the $k$ Gaussians that are closest to it to form the point set $M_{i_{K}}$. The local center of mass C of $M_{i_{K}}$ is given by the following equation:
@@
2.5 Structure Perception Optimization
@@
Loss. Combining image generation quality, human perception, and joint rotation constraints, we adopt the following Loss:
@@
2.5.1 Implementation Details.
We employ AdamW with different parameters for different modules. Experiments are conducted on a single RTX 3090 GPU. The model takes about 3,000 iterations or so and finally converges within 4 minutes (512 $\times$ 512 resolution). Furthermore, a comprehensive overview of our network structure and hyperparameters can be found in the appendix.
@@ 3.1 Experimental Setups
Comparison methods.
Compare MOSS with two categories of SOTA methods: Human NeRF-based methods, such as NeuralBody [36], HumanNeRF [47] AnimateNeRF [35], InstantNVR [10] InstantAvatar [17]. 3DGS methods, including GauHuman [15] and Human101 [25]. Following  [10], for a fair comparison, we do not optimize the test time for the SMPL parameters using images from the test set on InstantAvatar [17]. More details see the appendix.
Metrics. PSNR, SSIM, and LPIPS* are utilized as indicators of quality, while FPS is employed as a metric for speed. PSNR evaluates the variance between pixel values of two images. SSIM measures image similarity in terms of luminance, contrast, and structural integrity. LPIPS* compares two images in the feature space, which is considered to be closer to human visual perception.
@@ 3.2 Evaluation
Reconstruction Efficacy of MOSS. We validate the constraint effect of MOSS on 3D Gaussians, as well as its ability to identify local regions with significant surface deformations. As shown in Figure 5, we select the same human body from two different viewpoints and the qualitative comparisons. The details of the human body joints are marked by the red box, and the details of the clothing folds and other deformations are marked by the yellow box. It shows that our method can reproduce the surface details and clothing folds with high quality in large motion scenarios.
@@
Figure 5: To ensure a fair comparison, we compare  [36, 47, 35, 10, 15, 25] at 512 $\times$ 512 resolution. Our model shows better visual quality and more detail. More comparative results are detailed in the appendix.
Gaussian Guidance. As shown in Figure 5, MOSS achieves more excellent visual quality in the details of human joints. The edges of clothing and joints are clearly demarcated, and the orientation of clothing folds is closer to the real occasion (e.g., shorts and thighs). Among the compared methods, NeuralBody encodes human body structure with the latent code of SMPL vertices, but it cannot capture complex human movement. This results in blurred fists and incomplete hands in the figure. AnimateNeRF expects to overcome the ill-posed problem. The human joint skeleton is introduced as regularisation to facilitate training. Its rendering results still have artifacts in arms and calves. GauHuman’s Gaussians lack global constraints, leading to problem such as confusion between shorts and knees, as well as loss of hand details. The KGAS module of MOSS makes the human body reconstruction constrained by the topology of the kinematic tree and guides the 3D Gaussian displacements with the global motion, which is able to fully learn the motion of human, and the differences between the clothes and the human body. In addition, the orientation of clothing folds is also more realistic due to the constraints of motion direction.
Surface Deformation Detection. MOSS restores more surface details. On the contrary, InstantNVR’s modeling only considers the human body and ignores the clothing, resulting in a complete loss of clothing folds of the human reconstructed by InstantNVR. What’s more, the InstantNVR rendered shorts tightly clinging to the legs, which is different from the loose shorts in reality. The pixel-based loss used in GauHuman only provides weak supervision, which results in its restoration of only a few folds (e.g., the sides of the waistcoat and shorts on the torso shown in the yellow box in the first row of results). The UID detects complex deformations on the local body surface and compensates for the geometric reconstruction. Hence, UID effectively solves the problem of insufficient supervision in a single-view, and the folds that are hidden from each other can also be restored.
Table 2: We use Sequence 377 from ZJU-MoCap to test the effectiveness of Joint-Driven Orientation Refinement and its components.
@@
JNTM (Hierarchical + Joint Driver) 24.98 30.87 0.971
@@
Table 3: These evaluations are conducted using Sequence 377 from the ZJU-MoCap dataset, ensuring that each component of MOSS works.
@@ 3.3 Ablation Study
We conduct ablation studies to examine the role that every component plays in MOSS. The results of our experiments on the Sequence 377 from ZJU-MoCap are shown in Table 2, Table 3, Figure 6 and Figure 7. Table 2 presents the effectiveness of JNTM and its sub-components. The baseline is 3D Gaussian Splatting initialized with SMPL. By adding Autoregression and Joint Driver incrementally, the quality of human body reconstruction is progressively improved. Table 3 shows the results of MOSS and the outcomes after removing only one component each time. Through selectively disabling each component, we evaluate their contribution to the system. We will analyze the efficacy of individual components according to experimental results specifically in the following part.
@@
Joint-Driven Orientation Refinement. As is shown in Table 2, the Autoregression improves the performance in LPIPS*, PSNR, and SSIM. This validates that it is effective to use the kinematic tree to constrain the parameter estimation for the distribution of joint rotations. Furthermore, Joint Driver also improves the metrics, this is because it precisely controls Gaussians by joint movement. In comparison, Ours has significantly enhanced the LPIPS*, owing to the incorporation of guidance on Gaussians and surface deformation detection. The rendered and depth images in Figure 6 (a) and Figure 7 also show the progressively improved effect when sequentially adding subcomponents of JNTM.
@@
Density Perceptual Sampler. We conduct the following ablation experiments to explore the Gaussian density and surface structure perceptual ability. The results are shown in Table 3 w/o Density Perceptual Sampler. According to the green box in Figure 6 (b), the absence of this component results in the blurred rendering of surface edges, especially the fingers and the face. It indicates that the Gaussians of the human body and surface are severely underdense on the boundaries in Figure 7. In comparison, the Density Perceptual Sampler uses the principal axis concentration to correct the sampling probability of the Gaussian split process, such that the sampling area is more aligned with the actual surface. Besides, the restored deformation is more significant and clear even if the human is performing motions such as head rotation, finger clenching, lumbar rotation, and foot padding.
@@
Motion Guidance. The experimental results highlight the impact of Motion Guidance. Human body movements, such as fist clenching or arm raising, lead to local shadows. Comparison of blue boxes in Figure 6 (b) and Figure 7 shows that Motion Guidance can mitigate the confusion between shadows and actual body parts. Additionally, the blue circle illustrates that Motion Guidance helps to restore the precise shapes of irregular and flexible clothing (e.g., belt loop). This component guides Gaussians with global motion, dynamically adjusting the orientation and scaling of each Gaussian. Consequently, it not only enhances the fidelity of reproducing local shadows in human body motion but also improves the accuracy in describing texture and shape alterations in clothing. The efficacy of w/o Motion Guidance is further elucidated through the results of ablation experiments in Table 3.
@@
UID. To verify UID’s ability to identify the complex folds on the surface, we perform the following ablation experiments, with the results shown in Figure 6 (b). If we omit UID and only use pixel-wise losses for supervision, it will lead to insufficient guidance. Especially in Figure 7, most of the body parts lack depth information. Hence, it gets more difficult to accurately reconstruct various human body surfaces (especially under occlusion scenarios, seen in Figure 4). Furthermore, a comparison between the metrics of w/o UID and Full in Table 3 reveals a decline in the performance across all evaluation metrics. Confusion between the human body and trunk clothing (see yellow boxes), surface rendering distortions (see the yellow circle), and significant loss of facial features, including drastic texture changes and loss of information in high-frequency regions (see the yellow polygon), are evident. In addition, Figure 6 also shows that without UID leads to inaccurate restoration of overlapping lower legs, knees, and shorts with significant artifacts.

### v2

@@ 3.4 Surface Deformation Detector(UID)
We densify the Gaussians based on gradient localization to increase the rendering granularity for better reproduction of clothing folds and joint details. Remember that the set of Gaussians selected by the gradient is $M=\{m_{1},m_{2},...\}$. For each point $m_{i}$ in the point set $M$, select the $k$ Gaussians that are closest to it to form the point set $M_{i_{K}}$. The local center of mass C of $M_{i_{K}}$ is given by the following equation:
@@
3.5 Loss Function and Training Optimization
@@
3.5.1 Loss
Combining image generation quality, human perception, and joint rotation constraints, we adopt the following Loss:
@@ 4.1.1 Datasets
4.1.2 Comparison methods
Compare MOSS with two categories of SOTA methods: Human NeRF-based methods, such as NeuralBody [39], HumanNeRF [50] AnimateNeRF [38], InstantNVR [10] InstantAvatar [17]. 3DGS methods, including GauHuman [15] and Human101 [26]. Following [10], for a fair comparison, we do not optimize the test time for the SMPL parameters using images from the test set on InstantAvatar [17].
4.1.3 Metrics
PSNR, SSIM, and LPIPS* are utilized as indicators of quality, while FPS is employed as a metric for speed. PSNR evaluates the variance between pixel values of two images. SSIM measures image similarity in terms of luminance, contrast, and structural integrity. LPIPS* compares two images in the feature space, which is considered to be closer to human visual perception.
4.1.4 Implementation Details
We utilize AdamW with varying parameters across different modules. Experiments were carried out on a single RTX 3090 GPU, with the model taking approximately 3,000 iterations to converge. Additionally, a detailed overview of our network structure and hyperparameters can be found in the appendix.
@@ 4.2 Evaluation
Figure 5: To ensure a fair comparison, we compare [39, 50, 38, 10, 15, 26] at 512 $\times$ 512 resolution. Our model shows better visual quality and more detail.
4.2.1 Reconstruction Efficacy of MOSS
We validate the constraint effect of MOSS on 3D Gaussians, as well as its ability to identify local regions with significant surface deformations. As shown in Figure 5, we select the same human body from two different viewpoints and the qualitative comparisons. The details of the human body joints are marked by the red box, and the details of the clothing folds and other deformations are marked by the yellow box. It shows that our method can reproduce the surface details and clothing folds with high quality in large motion scenarios.
4.2.2 Gaussian Guidance
As shown in Figure 5, MOSS achieves more excellent visual quality in the details of human joints. The edges of clothing and joints are clearly demarcated, and the orientation of clothing folds is closer to the real occasion (e.g., shorts and thighs). Among the compared methods, NeuralBody encodes human body structure with the latent code of SMPL vertices, but it cannot capture complex human movement. This results in blurred fists and incomplete hands in the figure. AnimateNeRF expects to overcome the ill-posed problem. The human joint skeleton is introduced as regularisation to facilitate training. Its rendering results still have artifacts in arms and calves. GauHuman’s Gaussians lack global constraints, leading to problem such as confusion between shorts and knees, as well as loss of hand details. The KGAS module of MOSS makes the human body reconstruction constrained by the topology of the kinematic tree and guides the 3D Gaussian displacements with the global motion, which is able to fully learn the motion of human, and the differences between the clothes and the human body. In addition, the orientation of clothing folds is also more realistic due to the constraints of motion direction.
4.2.3 Surface Deformation Detection
MOSS restores more surface details. On the contrary, InstantNVR’s modeling only considers the human body and ignores the clothing, resulting in a complete loss of clothing folds of the human reconstructed by InstantNVR. What’s more, the InstantNVR rendered shorts tightly clinging to the legs, which is different from the loose shorts in reality. The pixel-based loss used in GauHuman only provides weak supervision, which results in its restoration of only a few folds (e.g., the sides of the waistcoat and shorts on the torso shown in the yellow box in the first row of results). The UID detects complex deformations on the local body surface and compensates for the geometric reconstruction. Hence, UID effectively solves the problem of insufficient supervision in a single-view, and the folds that are hidden from each other can also be restored.
Table 2: The effectiveness of Joint-Driven Orientation Refinement and its components
@@ 4.2.3 Surface Deformation Detection
JNTM (Hierarchical + joint linkage) 24.98 30.87 0.971
@@
Table 3: Evaluations to ensure the functionality of each component of MOSS.
@@ 4.3 Ablation Study
We conduct ablation studies to examine the role that every component plays in MOSS. The results of our experiments on ZJU-MoCap are shown in Table 5, Table 6, Figure 6 and Figure 7. Table 5 presents the effectiveness of JNTM and its sub-components. The baseline is 3D Gaussian Splatting initialized with SMPL. By adding Autoregression and joint linkage incrementally, the quality of human body reconstruction is progressively improved. Table 6 shows the results of MOSS and the outcomes after removing only one component each time. Through selectively disabling each component, we evaluate their contribution to the system. We will analyze the efficacy of individual components according to experimental results specifically in the following part.
4.3.1 Joint-Driven Orientation Refinement
@@
As is shown in Table 5, the Autoregression improves the performance in LPIPS*, PSNR, and SSIM. This validates that it is effective to use the kinematic tree to constrain the parameter estimation for the distribution of joint rotations. Furthermore, joint linkage also improves the metrics, this is because it precisely controls Gaussians by joint movement. In comparison, Ours has significantly enhanced the LPIPS*, owing to the incorporation of guidance on Gaussians and surface deformation detection. The rendered and depth images in Figure 6 (a) and Figure 7 also show the progressively improved effect when sequentially adding subcomponents of JNTM.
4.3.2 Density Perceptual Sampler
@@
We conduct the following ablation experiments to explore the Gaussian density and surface structure perceptual ability. The results are shown in Table 6 w/o Density Perceptual Sampler. According to the green box in Figure 6 (b), the absence of this component results in the blurred rendering of surface edges, especially the fingers and the face. It indicates that the Gaussians of the human body and surface are severely underdense on the boundaries in Figure 7. In comparison, the Density Perceptual Sampler uses the principal axis concentration to correct the sampling probability of the Gaussian split process, such that the sampling area is more aligned with the actual surface. Besides, the restored deformation is more significant and clear even if the human is performing motions such as head rotation, finger clenching, lumbar rotation, and foot padding.
4.3.3 Motion Guidance
@@
The experimental results highlight the impact of Motion Guidance. Human body movements, such as fist clenching or arm raising, lead to local shadows. Comparison of blue boxes in Figure 6 (b) and Figure 7 shows that Motion Guidance can mitigate the confusion between shadows and actual body parts. Additionally, the blue circle illustrates that Motion Guidance helps to restore the precise shapes of irregular and flexible clothing (e.g., belt loop). This component guides Gaussians with global motion, dynamically adjusting the orientation and scaling of each Gaussian. Consequently, it not only enhances the fidelity of reproducing local shadows in human body motion but also improves the accuracy in describing texture and shape alterations in clothing. The efficacy of w/o Motion Guidance is further elucidated through the results of ablation experiments in Table 6.
4.3.4 UID
@@
To verify UID’s ability to identify the complex folds on the surface, we perform the following ablation experiments, with the results shown in Figure 6 (b). If we omit UID and only use pixel-wise losses for supervision, it will lead to insufficient guidance. Especially in Figure 7, most of the body parts lack depth information. Hence, it gets more difficult to accurately reconstruct various human body surfaces (especially under occlusion scenarios, seen in Figure 4). Furthermore, a comparison between the metrics of w/o UID and Full in Table 6 reveals a decline in the performance across all evaluation metrics. Confusion between the human body and trunk clothing (see yellow boxes), surface rendering distortions (see the yellow circle), and significant loss of facial features, including drastic texture changes and loss of information in high-frequency regions (see the yellow polygon), are evident. In addition, Figure 6 also shows that without UID leads to inaccurate restoration of overlapping lower legs, knees, and shorts with significant artifacts.